OpenStudy (anonymous):

you need to purchase a textbook for nursing school. the book cost 80.00 and the sales tax is 8.25 where you are purchasing the book if you have $100 how much change will you receive back

OpenStudy (amistre64):

is the tax $8.25? or is that supposed to be: 8.25% ??

OpenStudy (anonymous):

yes the tax is 8.25

OpenStudy (amistre64):

ill assume its (%) 80 + 80(decimal of percent) = total cost: $100 - total cost = change back. 80 + 80(.0825) = Tc 80(1+.0825) = Tc 80(1.0825) = Tc 100 - 80(1.0825) = change back 13.40 is what youll get back if the tax is 8.25%

OpenStudy (anonymous):

ok hold on let me process all of it lol

OpenStudy (anonymous):

ok so when you say $100.00- total cost does that look like this 100.00-80.00

OpenStudy (amistre64):

it looks kind of like that. We can look at it this way to: $100 - (book price) - (taxes paid on book price) The $100 is money you used to pay for the book; so we are expecting change back for it right?

OpenStudy (amistre64):

100 - 80 - (taxes) = change gotten back

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ooo ok let me work it out on my own and ill let you know what i get thank you :D

OpenStudy (anonymous):

80 x 1.0825-100=13.40

OpenStudy (amistre64):

i gotta head to marketing class now; and study for a macroeconomics exam afterwards.... Ciaobella :)

OpenStudy (anonymous):

you just subtract the total cost out of the $100 ( I am assuming a tax of $8.25 in each book).. 100-the total cost= change back... the total cost=80+8.25=88.25 therefor the change back=100-88.25=11.75

OpenStudy (anonymous):

simply: 100-(80+8.25)=11.75